# The First and Second-Order Asymptotics of Covert Communication over AWGN Channels

### v5

@@ Abstract
This paper establishes the first- and second-order asymptotics of the maximal throughput under a covertness criterion defined by a Kullback-Leibler (KL) divergence bound $\delta$. The achievability employs random coding with three key innovations: (1) dynamically scaled power constraints dependent on blocklength $n$, (2) codeword generation via truncated Gaussian distributions optimized for near-isotropy, and (3) codebook expurgation to ensure reliability. The analysis reveals two critical results. First, the maximal throughput scales as $O(n^{\frac{1}{2}})$ with second-order term $O(n^{\frac{1}{4}})$. Compared with the previous results, we present analytic formulas of the asymptotics as functions of the truncation hyper-parameter $\mu$, the adversary’s noise variance $\sigma^{2}$ and the upper bound $\delta$ on KL divergence. Second, the required secret key size exhibits a universal scaling law of $O(n^{\frac{1}{2}})$ even when the legitimate receiver’s channel quality marginally exceeds the adversary’s—a sharp contrast to discrete memoryless channels where such asymmetry reduces key requirements. Theoretical contributions are further reinforced through an information-geometric perspective: truncated Gaussians minimize detectability by confining power within $O(n^{\frac{1}{2}})$ while maximizing throughput through near-capacity-achieving properties. The square-root power law is further justified via asymptotic analyses of several divergences of Gaussian family with different variances. These results advance covert communication design by providing explicit scaling laws, optimal codebook architectures, and a systematic methodology for balancing reliability, covertness, and resource constraints.
@@ I Introduction
Considering the latency and practical length of the codewords, the finite blocklength performance [20] is important for practical performance evaluation. The characterization on the second order asymptotics of the maximal throughput over covert channels is necessary in a finite blocklength analysis, which represents the backoff quantity with given finite blocklength $n$, probability $\epsilon$ of decoding error and some covert metric $\delta$ from the first order asymptotics. Tahmasbi and Bloch characterize the second order asymptotics of covert communication over discrete memoryless channels in [21][22] from the perspective of channel resolvability. A closed form of the second order asymptotics is given when the covert metric is KL divergence, and related bounds are provided when the covert metrics are total variation distance (TVD) and missed detection probability, respectively. In [7] and [23], one-shot achievability bound of Gaussian random coding over AWGN channels under maximal power constraint are presented together with the TVD between the distributions of channel outputs at the adversary with and without communications of the legitimate transmitter. The converse bound under a maximal power constraint that is varying with blocklength $n$ is also provided. Nevertheless, the characterization of the first and second order asymptotics as a function of the covert constraint $\delta$ over AWGN channels is still missing.
@@
This paper investigates the first- and second-order asymptotics of covert communication over AWGN channels, characterized as functions of blocklength $n$ decoding error probability $\epsilon$, the noise level at the adversary $\sigma$ and covertness parameter $\delta$, herein $\delta$ constrains the Kullback-Leibler (KL) divergence between the adversary’s observation distributions with and without communication. For achievability analysis, we extend our previous framework in [23] by maintaining the truncated Gaussian generating distribution with hyper-parameter $\mu$, ̵while replacing the sequential encoding scheme with an enhanced approach combining i.i.d. sampling and expurgation operations from [22]. Our reliability analysis introduces a novel decoding threshold selection mechanism that effectively bounds error probabilities for codewords with varying radii – ̵a distinctive feature not addressed in prior works [20, 24]. The channel resolvability analysis advances in two key aspects: First, ̵we circumvent the quasi-metric limitations of KL divergence in Euclidean space by directly analyzing the divergence between the Gaussian mixture induced by codewords and the noise distribution, avoiding reliance on modified triangle inequalities applicable to DMCs. Second, ̵we establish an explicit connection between adversary noise levels and required key bits, revealing a fundamental dichotomy between AWGN and DMC scenarios. ̵Notably, the key size remains $O(\sqrt{n})$ even when adversary noise marginally exceeds main channel noise – a critical divergence from DMC behavior that we rigorously explain through channel characteristic analysis. The optimality of truncated Gaussian distributions is justified through local information geometry principles [26, 27], where covertness realization corresponds to positioning the signal distribution within a quasi-$\varepsilon$ neighborhood of noise distribution in statistical manifold space. ̵This geometric interpretation bridges coding theory with machine learning applications [28, 29]. The square root law governing power scaling $\Psi(n)=O(\frac{1}{\sqrt{n}})$ emerges naturally from asymptotic analyses of Gaussian divergence measures and noise distribution approximation accuracy. Note that our work only considers AWGN channel model and do not take into account of the signal duration or spectral width. Under such circumstance, the transmission involves specific modulation, and has been investigated in [30][31].
@@
The rest of this paper is organized as follows. In Section II, we present some preliminaries on notations, the channel model, the outline of the proof of the achievability and the converse, the definitions of codes and some distributions. Section III presents the main results, which include the first and second order asymptotics of the achievability bound and the investigation of the converse bound. The framework of local information geometry to explain the optimality of truncated Gaussian distribution is provided in Section IV. The investigation on the connection between Square Root Law, the choice of the power level and the asymptotic behavior of several divergences of Gaussian distributions is provided in Section V. Section VI concludes the paper.
@@ II-C Definitions for Codes over AWGN Covert Channels
A $(M,n,\epsilon,\Psi)_{m}$ code $\mathcal{C}=\{\bm{c}_{1},\bm{c}_{2},\cdots,\bm{c}_{M}\}$ for a AWGN channel (5), consists of a message set $W\in\langle M\rangle$, an encoder at the transmitter Alice $f_{n}:\mathcal{W}\rightarrow\mathbb{R}^{n},w\mapsto\bm{x}$, and a decoder at the legitimate user Bob $g_{n}:\mathbb{R}^{n}\rightarrow\langle M\rangle,\bm{y}\rightarrow\hat{w}$ with the condition that the encoder and decoder pair satisfies
@@
    maximal power constraint: each codeword $\bm{c}_{i}$ satisfies
    $\|\bm{c}_{i}\|^{2}\leq n\Psi,$ (8)
@@
    $\frac{1}{M}\sum_{w=1}^{M}\|\bm{x}_{sw}\|^{2}\leq n\Psi\,\ \,\text{holds for each}\,\ \,w\in\langle M\rangle$ (15)
The maximal sizes of $(M,n,\epsilon,\Psi)_{m}$, $(M,n,\epsilon,\Psi)_{a}$, $(M,K,n,\epsilon,\Psi,\sigma,\delta)_{m}$, $(M,K,n,\epsilon,\Psi,\sigma,\delta)_{a}$, $(M,K,n,\epsilon,\sigma,\delta)$ codes are denoted by $M^{*}_{m}(n,\epsilon,\Psi)$, $M^{*}_{a}(n,\epsilon,\Psi)$, $M^{*}_{m}(n,\epsilon,\Psi,\sigma,\delta)$, $M^{*}_{a}(n,\epsilon,\Psi,\sigma,\delta)$ and $M^{*}(n,\epsilon,\sigma,\delta)$ respectively. In addition, for any $(M,K,n,\epsilon,\sigma,\delta)$ code, the corresponding number of key bits is denoted as $K(n,\epsilon,\sigma,\delta)$.
@@
As the distribution $Q^{(n)}_{\bm{Z},\bm{C}}(\bm{z})$ depends on the codewords $\bm{x}_{sw}$ and the codewords are randomly generated, the quantity $Q^{(n)}_{\bm{Z},\bm{C}}(\bm{z})$ itself is a random variable.
@@
II-D Outline of The Proof
In this section, we provide a road map on the proof of the achievability and converse bounds.
There are mainly four steps for achievability.
@@ II-D Outline of The Proof
The proof of the converse consists of two steps.
@@
    By using the pigeon-hole principle, the converse bound for each sub-code under maximal power constraint is derived in Corollary 3, which further leads to characterization of the converse bound under covert constraint in Theorem 7.
@@ II-E Related Distributions
The proof can be found in Appendix B. With $0<\mu<1$ being a constant, $\Delta\rightarrow 1$ as $n\rightarrow\infty$ exponentially and the rapidity depends only on $\mu$. Please see Fig.2 for reference. Thus, we have the following corollary for $\mathbb{V}(\bar{P}_{n}^{(n)},P_{n}^{(n)})$, the details of the analysis has been presented in Section IV-A of [23].
Fig. 2: $\Delta$ tends to 11 with proper chosen $\mu$.
@@
The distribution pair $\bar{P}_{n}^{(n)}$ and $P_{n}^{(n)}$ satisfies
@@
For sufficiently large $n$, for any fixed $0<\mu<1$ and any $\bm{z}$, we have
@@ III-A Random Coding, Decoding and Reliability Analysis
$\mathsf{V}(\mu\Psi(n)):=\mathsf{V}(\mu\Psi(n),\mu\Psi(n))=\frac{2\mu\Psi(n)+\mu^{2}\Psi^{2}(n)}{2(1+\mu\Psi(n))^{2}}.$ (56)
@@
$\overset{(c)}{\leq}$ $\Phi\left(\Phi^{-1}\big(\epsilon_{1}-\frac{\underset{\Upsilon\in[\mu^{2}\Psi(n),\Psi(n)]}{\max}B(\Upsilon,\mu\Psi(n))}{\sqrt{n}}\big)\right)+\frac{B(\Upsilon,\mu\Psi(n))}{\sqrt{n}}$
@@
In the $\kappa\beta$ bound (Theorem 25) of [20], the codewords lie on the $n$-dimensional sphere $\mathcal{S}^{n-1}$ with radius $\sqrt{nP}$ and are not randomly generated by any distribution. Nevertheless, the auxiliary distribution $\mathcal{N}(0,(1+P)\bm{I}_{n})$ is chosen so that normal approximation can be performed. In this work, the generating distribution $\bar{P}_{n}^{(n)}$ is chosen such that its induced output distribution $\bar{Q}^{(n)}_{\bm{Y},1}$ is close to the auxiliary distribution $Q^{(n)}_{\bm{Y},1}$, which makes the normal approximation accurate. The bounding of the probability of error $\underset{s,w}{\max}\,\ \,\varepsilon_{sw}^{(1)}$ relies on the fact that the $l^{2}$ norms of the codewords lie in a small interval $\mu^{2}n\Psi(n),n\Psi(n)]$. Our strategy is separately consider $n$-dimensional sphere $\mathcal{S}^{n-1}$ with radius from $\sqrt{\mu^{2}n\Psi(n)}$ to $\sqrt{n\Psi(n)}$, and for each sphere, we can utilize the sphere symmetry to simplify the analysis. The parameter $\mu$ makes the interval under control so that an universal decoding threshold $\gamma$ can be obtained. As a consequence, there is an additional term $\bar{\mathsf{C}}(\Upsilon,\mu\Psi(n))$ in the normal approximation, which is different from the previous works [24][20] where all the codewords have the same power level.
@@
In our previous work [23], the encoder adopts sequential random coding (Theorem 1 in [23]) so that the codewords are not i.i.d generated, which makes it difficult to perform resolvability analysis. In this work, the codewords are i.i.d sampled from the truncated Gaussian distribution $\bar{P}_{n}^{(n)}$ that is close to Gaussian family. The resolvability analysis on bounding the induced output distribution of the codewords and the distribution of noise will depend much on this fact.
Remark 7.
In RCU bound in [24], the codewords are uniformly distributed on $\mathcal{S}^{n-1}$, and maximal likelihood decoding is adopted. The evaluation of it relies on the high-probability set $\mathcal{F}$ (Please refer to (31) in Section IV-D of [24]) and the fact that the induced output distribution and the auxiliary distribution can be bounded by a finite constant:
@@
Note that for the error probability of the second type $\varepsilon_{sw}^{(2)}$, we can not bound it for each codeword. Nevertheless, our strategy is divided into two steps. First, we use an existing results to bound the average of it in Lemma 6. Then, we leverage the expurgation later to discard part of codewords with large error probability of the second type. The following lemma is almost a restatement of part of Lemma 1 in [22]. The differences are: (1) The discrete memoryless channel is substituted by the AWGN channel. (2) The sum is substituted by integration. (3) We use $\log$ with base $2$ instead of base $e$.
@@
$\times$ $\Phi^{-1}(\epsilon_{1}-\frac{\underset{\Upsilon\in[\mu^{2}\Psi(n),\Psi(n)]}{\max}B(\Upsilon,\mu\Psi(n))}{\sqrt{n}})$
@@
The proof is obvious from (63) and (4). Note that here we just provide an upper bound of the average error probability of the second type $\varepsilon_{sw}^{(2)}$. To get a reliable code, we will adopt expurgation later in Theorem 6 to construct a reliable code such that the maximum of the sum $\varepsilon_{sw}^{(1)}+\varepsilon_{sw}^{(2)}$ for all $(s,w)$ pair is small.
@@
From (65), the average of the decoding error probability of the second type depends on two elements. The first element is $\mathbb{E}_{\bar{Q}^{(n)}_{\bm{Y},1}}\bigg(\frac{\bar{g}_{1}^{(n)}(\bm{y})}{g_{1}^{(n)}(\bm{y})}\bigg)$. The integrand is the ratio of the probability density functions of the induced output distribution of the generating distribution and the reference distribution, where the latter is the induced distribution of Gaussian distribution without truncation. The second element is the size of the codebook, which has been almost determined in Lemma 4. Hence, it is necessary to control the effect of truncation to bound the ratio $\frac{\bar{g}_{1}^{(n)}(\bm{y})}{g_{1}^{(n)}(\bm{y})}$, which depends on the choice of the parameter $\mu$ and has been bounded for sufficiently large $n$ in Lemma 3.

### v6

@@ Abstract
This paper establishes the first- and second-order asymptotics of the maximal throughput under a covertness criterion defined by a Kullback-Leibler (KL) divergence bound $\delta$. The achievability employs random coding with three key innovations: (1) dynamically scaled power constraints dependent on blocklength $n$, (2) codeword generation via truncated Gaussian distributions optimized for near-isotropy, and (3) codebook expurgation to ensure reliability. The analysis reveals two critical results. First, compared with the previous results, we present analytic formulas of the first-and second order asymptotics as functions of blocklength $n$, decoding error probability $\epsilon$, the adversary’s noise variance $\sigma^{2}$ and the upper bound $\delta$ on KL divergence. Second, the required secret key size exhibits a universal scaling law of $O(n^{\frac{1}{2}})$ even when the legitimate receiver’s channel quality marginally exceeds the adversary’s — a sharp contrast to discrete memoryless channels where such asymmetry reduces key requirements. Theoretical contributions are further reinforced through an information-geometric perspective: truncated Gaussians minimize detectability by confining power within $O(n^{\frac{1}{2}})$ while maximizing throughput through near-capacity-achieving properties. The square-root power law is further justified via asymptotic analyses of several divergences of Gaussian family with different variances. These results advance covert communication design by providing explicit scaling laws, optimal codebook architectures, and a systematic methodology for balancing reliability, covertness, and resource constraints.
@@ I Introduction
Considering the latency and practical length of the codewords, the finite blocklength performance [20] is important for practical performance evaluation. The characterization on the second order asymptotics of the maximal throughput over covert channels is necessary in a finite blocklength analysis, which represents the back-off quantity with given finite blocklength $n$, probability $\epsilon$ of decoding error and some covert metric $\delta$ from the first order asymptotics. Tahmasbi and Bloch characterize the second order asymptotics of covert communication over discrete memoryless channels in [21][22] from the perspective of channel resolvability. A closed form of the second order asymptotics is given when the covert metric is KL divergence, and related bounds are provided when the covert metrics are total variation distance (TVD) and missed detection probability, respectively. In [7] and [23], one-shot achievability bound of Gaussian random coding over AWGN channels under maximal power constraint are presented together with the TVD between the distributions of channel outputs at the adversary with and without communications of the legitimate transmitter. The converse bound under a maximal power constraint that is varying with blocklength $n$ is also provided. Nevertheless, the characterization of the first and second order asymptotics as a function of the covert constraint $\delta$ over AWGN channels is still missing.
@@
This paper investigates the first- and second-order asymptotics of covert communication over AWGN channels, characterized as functions of blocklength $n$ decoding error probability $\epsilon$, the noise level at the adversary $\sigma$ and covertness parameter $\delta$, herein $\delta$ constrains the Kullback-Leibler (KL) divergence between the adversary’s observation distributions with and without communication. For achievability analysis, we extend our previous framework in [23] by maintaining the truncated Gaussian generating distribution with hyper-parameter $\mu$, ̵while replacing the sequential encoding scheme with an enhanced approach combining i.i.d. sampling and expurgation operations from [22]. Our reliability analysis introduces a novel decoding threshold selection mechanism that effectively bounds error probabilities for codewords with varying radii – ̵a distinctive feature not addressed in prior works [20, 24]. The channel resolvability analysis advances in two key aspects: First, ̵we circumvent the quasi-metric limitations of KL divergence in Euclidean space by directly analyzing the divergence between the Gaussian mixture induced by codewords and the noise distribution, avoiding reliance on modified triangle inequalities applicable to DMCs. Second, ̵we establish an explicit connection between adversary noise levels and required key bits, revealing a fundamental dichotomy between AWGN and DMC scenarios. ̵Notably, the key size remains $O(\sqrt{n})$ even when adversary noise marginally exceeds main channel noise – a critical divergence from DMC behavior that we rigorously explain through channel characteristic analysis. The optimality of truncated Gaussian distributions is justified through local information geometry principles [26, 27], where covertness realization corresponds to positioning the signal distribution within a quasi-$\varepsilon$ neighborhood of noise distribution in statistical manifold space. ̵This geometric interpretation bridges coding theory with machine learning applications [28, 29]. The square root law governing power scaling $\Psi(n)=O(\frac{1}{\sqrt{n}})$ emerges naturally from asymptotic analyses of Gaussian divergence measures and noise distribution approximation accuracy. Note that our work only considers AWGN channel model and its performance limit under finite blocklegnth regime. We do not take into account of modulation, the signal duration, or spectral width. Under such circumstance, the transmission involves specific modulation, and has been investigated in [30][31].
@@
The rest of this paper is organized as follows. In Section II, we present some preliminaries on notations, the channel model, the outline of the proof of the achievability and the converse, the definitions of codes and some distributions. Section III presents the main results, which include the first and second order asymptotics of the achievability bound and the investigation of the converse bound. The framework of local information geometry to explain the optimality of truncated Gaussian distribution, the investigation on the connection between Square Root Law, the choice of the power level and the asymptotic behavior of several divergences of Gaussian distributions are provided in Section IV. Section V concludes the paper.
@@ II-C Definitions for Codes over AWGN Covert Channels
A $(M,n,\epsilon,\Psi)_{m}$ code $\mathcal{C}=\{\bm{x}_{1},\bm{x}_{2},\cdots,\bm{x}_{M}\}$ for a AWGN channel (5), consists of a message set $W\in\langle M\rangle$, an encoder at the transmitter Alice $f_{n}:\mathcal{W}\rightarrow\mathbb{R}^{n},w\mapsto\bm{x}$, and a decoder at the legitimate receiver Bob $g_{n}:\mathbb{R}^{n}\rightarrow\langle M\rangle,\bm{y}\rightarrow\hat{w}$ with the condition that the encoder and decoder pair satisfies
@@
    maximal power constraint: each codeword $\bm{x}_{i}$ satisfies
    $\|\bm{x}_{i}\|^{2}\leq n\Psi,$ (8)
@@
    $\frac{1}{M}\sum_{w=1}^{M}\|\bm{x}_{sw}\|^{2}\leq n\Psi.$ (15)
The maximal sizes of $(M,n,\epsilon,\Psi)_{m}$, $(M,n,\epsilon,\Psi)_{a}$, $(M,K,n,\epsilon,\Psi,\sigma,\delta)_{m}$, $(M,K,n,\epsilon,\Psi,\sigma,\delta)_{a}$, $(M,K,n,\epsilon,\sigma,\delta)$ codes are denoted by $M^{*}_{m}(n,\epsilon,\Psi)$, $M^{*}_{a}(n,\epsilon,\Psi)$, $M^{*}_{m}(n,\epsilon,\Psi,\sigma,\delta)$, $M^{*}_{a}(n,\epsilon,\Psi,\sigma,\delta)$ and $M^{*}(n,\epsilon,\sigma,\delta)$ respectively. For AWGN covert channels, we are interested in the quantity $M^{*}(n,\epsilon,\sigma,\delta)$ as a function of $n,\epsilon,\sigma,\delta$ since the covert constraint (12) always implies power constraint on codewords. In addition, for any $(M,K,n,\epsilon,\sigma,\delta)$ code, we want to characterize the corresponding number of key bits, which is denoted as $K(n,\epsilon,\sigma,\delta)$.
@@
As the distribution $Q^{(n)}_{\bm{Z},\bm{C}}(\bm{z})$ depends on the codewords $\bm{x}_{sw}$, which are randomly generated, the quantity $Q^{(n)}_{\bm{Z},\bm{C}}(\bm{z})$ itself is a random variable.
@@
II-D Main Results and Outline of The Proof
In this section, we provide the main results and a road map on the proof of the achievability and converse bounds.
Theorem 1.
For AWGN covert channels with channel models in (5) and (6), we have
$+$ $\sigma(2\mu)^{1/2}(n\delta)^{1/4}(\log e)^{3/4}\Phi^{-1}(\epsilon)+o(n^{\frac{1}{4}}).$ (19)
and
$\log$ $M^{*}(n,\epsilon,\sigma,\delta)\leq\sigma^{2}(n\delta\log e)^{1/2}$
$+$ $\sigma(2)^{1/2}(n\delta)^{1/4}(\log e)^{3/4}\cdot\Phi^{-1}(\epsilon)+O(\log n).$ (20)
The corresponding number of key bits should satisfy
In (1) and (21), $\mu\in(\frac{1}{2},1)$ is a hyper-parameter.
There are mainly four steps to arrive at achievability in Theorem 1.
@@ II-D Main Results and Outline of The Proof
There are two steps to arrive at converse in Theorem 1.
@@
    By using the pigeon-hole principle, the converse bound for each sub-code under maximal power constraint is derived in Corollary 3, which further leads to characterization of the converse bound in Theorem 1.
For covert communication over AWGN channels, previous works [10] [11] have established the characterization of the first order asymptotics of the throughput as $O(\sqrt{n})$. This work focuses on refining these results by characterizing both the first and second order asymptotics as functions of the blocklength $n$, decoding error probability $\epsilon$, the background noise level parameter $\sigma$ and covert parameter $\delta$, which is not reported in [10] [11]. The analytic investigation stresses the influence of the ratio of the noise level (which is $\sigma^{2}/1$ in this work) on the throughput and the size of key bits. From Theorem 1, we can see that both of the first order asymptotics are $O(n^{\frac{1}{2}})$ and both of the second order asymptotic are $O(n^{\frac{1}{4}})$, and the only difference on the analytic coefficients is the hyper-parameter $\mu$ of truncation in the achievability. The truncation parameter $\mu$ lies in both the size of codebooks (1) and the size of the key (21), which is different from the analysis of the first order asymptotics of covert communication over AWGN channels [10] [11] [12]. The detailed explanation on our results will be provided in Section III and Section IV.
@@ II-E Related Distributions
The proof can be found in Appendix B. With $0<\mu<1$ being a constant, $\Delta\rightarrow 1$ as $n\rightarrow\infty$ exponentially and the rapidity depends only on $\mu$. Please see Fig.2 for reference. Intuitively, the truncation of the tail of the Gaussian family depends on the hyper-parameter $\mu$. When $\mu$ is close to $1$, the tail probability is exponentially decreasing. Thus, we have the following corollary for $\mathbb{V}(\bar{P}_{n}^{(n)},P_{n}^{(n)})$, the details of the analysis has been presented in Section IV-A of [23].
Fig. 2: $\Delta$ tends to 11 with proper chosen $\mu$, which is the numerical results from (32).
@@
For proper fixed $\mu$, the distribution pair $\bar{P}_{n}^{(n)}$ and $P_{n}^{(n)}$ satisfies
@@
For sufficiently large $n$, for any proper chosen $\mu>0.5$ and any $\bm{z}$, we have
@@ III-A Random Coding, Decoding and Reliability Analysis
$\mathsf{V}(\mu\Psi(n)):=$ $\mathsf{V}(\mu\Psi(n),\mu\Psi(n))$
$=$ $\frac{2\mu\Psi(n)+\mu^{2}\Psi^{2}(n)}{2(1+\mu\Psi(n))^{2}}.$ (59)
@@
$\overset{(c)}{\leq}$ $\Phi\left(\Phi^{-1}\bigg(\epsilon_{1}-\frac{\underset{\Upsilon\in[\mu^{2}\Psi(n),\Psi(n)]}{\max}B(\Upsilon,\mu\Psi(n))}{\sqrt{n}}\bigg)\right)+\frac{B(\Upsilon,\mu\Psi(n))}{\sqrt{n}}$
@@
In the $\kappa\beta$ bound (Theorem 25) of [20], the codewords lie on the $n$-dimensional sphere $\mathcal{S}^{n-1}$ with radius $\sqrt{nP}$ and are not randomly generated by any distribution. Nevertheless, the auxiliary distribution $\mathcal{N}(0,(1+P)\bm{I}_{n})$ is chosen so that normal approximation can be performed. In this work, the codewords lie on the gray region as in Fig. 3.
Fig. 3: The codewords lie in the gray region.
The generating distribution $\bar{P}_{n}^{(n)}$ is chosen such that its induced output distribution $\bar{Q}^{(n)}_{\bm{Y},1}$ is close to the auxiliary distribution $Q^{(n)}_{\bm{Y},1}$, which makes the normal approximation accurate. The strategy of bounding the error probability of the first type here is separately considering $n$-dimensional sphere $\mathcal{S}^{n-1}$ with radius $\sqrt{n\Upsilon}$, which ranges from $\sqrt{\mu^{2}n\Psi(n)}$ to $\sqrt{n\Psi(n)}$. For each sphere, we can utilize the sphere symmetry to simplify the analysis. The parameter $\mu$ makes the interval under control so that an universal decoding threshold $\gamma$ can be obtained. As a consequence, there is an additional term $\bar{\mathsf{C}}(\Upsilon,\mu\Psi(n))$ in the normal approximation, which is different from the previous works [24][20] where all the codewords have the same power level. The bounding of the probability of error $\underset{s,w}{\max}\,\ \,\varepsilon_{sw}^{(1)}$ relies on the following facts. As $\mu$ is close to $1$, the $l^{2}$ norms of the codewords lie in a small interval $\mu^{2}n\Psi(n),n\Psi(n)]$. If we choose a small $\mu$ (close to 0), the size of the codebook should also be small so as to bound the error probability of the first type. This will not emerge when we only consider the first order asymptotics.
@@
In the evaluation of RCU bound in [24], the codewords are uniformly distributed on $\mathcal{S}^{n-1}$ with radius $\sqrt{nP}$, and maximal likelihood decoding is adopted. The derivation relies on the high-probability set $\mathcal{F}$ (Please refer to (31) in Section IV-D of [24]) and the fact that the induced output distribution and the auxiliary distribution can be bounded by a finite constant:
@@
The last lemma has provided a bound on error probability of the first type for all candidate codewords as: $\underset{s,w}{\max}\,\ \,\varepsilon_{sw}^{(1)}\leq\epsilon_{1}$ for given $\epsilon_{1}$. While, for the error probability of the second type $\varepsilon_{sw}^{(2)}$, we can not bound it for each codeword. Nevertheless, our strategy is divided into two steps. First, we use an existing results to bound the average of it in Lemma 6. Then, we leverage the expurgation later to discard part of codewords with large error probability of the second type. The following lemma is almost a restatement of part of Lemma 1 in [22]. The differences are: (1) The discrete memoryless channel is substituted by the AWGN channel. (2) The sum is substituted by integration. (3) We use $\log$ with base $2$ instead of base $e$.
@@
$\times$ $\Phi^{-1}(\epsilon_{1}-\frac{\max_{\Upsilon}B(\Upsilon,\mu\Psi(n))}{\sqrt{n}})$
@@
The proof is obvious from (66) and (4). Note that here we just provide an upper bound of the average error probability of the second type $\varepsilon_{sw}^{(2)}$. To get a reliable code, we will adopt expurgation later in Theorem 7 to construct a reliable code such that the maximum of the sum $\varepsilon_{sw}^{(1)}+\varepsilon_{sw}^{(2)}$ for all $(s,w)$ pairs is small.
@@
From (68), the average of the decoding error probability of the second type depends on two elements. The first element is $\mathbb{E}_{\bar{Q}^{(n)}_{\bm{Y},1}}\bigg(\frac{\bar{g}_{1}^{(n)}(\bm{y})}{g_{1}^{(n)}(\bm{y})}\bigg)$. The integrand is the ratio of the probability density functions of the induced output distribution of the generating distribution and the reference distribution, where the latter is the induced distribution of Gaussian distribution without truncation. The second element is the size of the codebook, which has been almost determined in Lemma 4. Hence, it is necessary to control the effect of truncation to bound the ratio $\frac{\bar{g}_{1}^{(n)}(\bm{y})}{g_{1}^{(n)}(\bm{y})}$, which depends on the choice of the parameter $\mu$ and has been uniformly bounded for sufficiently large $n$ in Lemma 3.